# Fairness Evaluation of Large Language Models in Academic Library Reference Services

### v2

@@ 5.3.1 General Behaviors
Beyond domain-specific vocabulary, LLMs also modulate service-oriented language based on assumptions about user autonomy and institutional familiarity. The use of “librarian” represents self-referential professional identification, as the LLM identifies itself as the service provider, and presents a particularly interesting case of professional authority assertion. This term appears more frequently in responses to outside users (+4.3) and alumni (+2.5), functioning as a form of institutional self-identification. This pattern may suggest that LLMs strategically deploy professional identity markers when interacting with patron types who are more likely to respond to explicit signals of institutional expertise and legitimacy as a means of establishing credibility (Drew and Heritage,, 1992).
The term “help” shows strong positive associations with outside users (+8.5) and faculty (+4.6), suggesting explicit assistance framing that serves different functions: providing institutional guidance for outside users while offering specialized research support for faculty. Conversely, “access” demonstrates consistent negative associations across most patron types, with particularly strong negative coefficients for graduate students (--8.1), alumni (--7.8), and outside users (--7.8). This pattern suggests that LLMs assume these groups either have existing access privileges or require different types of resource provision than undergraduates. The term “questions” shows negative associations with most patron types; these are particularly strong for graduate students (--11.1) and alumni (--9.1). This potentially indicates reduced meta-discourse about inquiry processes for users assumed to have greater institutional sophistication.
5.3.2 LLM-Specific Behaviors
@@
While the aggregate patterns demonstrate consistent role-based accommodation across LLMs, individual LLMs exhibit variable behaviors in implementing those accommodations. To illustrate this cross-model heterogeneity, we focus on three representative consensus features: “thank,” “dear,” and “research.” Figure 4 presents radar plots displaying each LLM’s coefficients for these features across patron types (see Appendix D for details).
@@
The “thank” panel reveals substantial variation in accommodation strategies across models. Three models (Gemma-2, GPT-4o, and Claude-3.5) exhibit zero coefficients across all patron types, indicating uniform usage regardless of user identity. In contrast, two models implement pronounced hierarchical accommodation: Llama-3.1 demonstrates positive associations across all patron types (staff +39.3, faculty +31.8, graduate students +25.3), while Gemini-2.5 shows similar positive patterns with different magnitudes (graduate students +32.0, staff +28.9, alumni +21.3). Ministral exhibits a distinct pattern with consistent negative coefficients across all groups (--5.9 to --12.0), suggesting reduced usage of “thank” for non-undergraduate patrons.
@@
The “dear” panel illustrates substantial model-specific variation in formality management approaches. GPT-4o exhibits the most pronounced pattern: positive coefficients for graduate students (+53.6) and faculty (+44.9) but negative coefficients for staff (--28.7) and outside users (--35.9), implementing a selective graduate/faculty-focused formality strategy. Gemini-2.5 is most formal toward undergraduate students. It otherwise demonstrates consistent formality reduction with negative coefficients across most patron types (--51.8 to --60.9), except for relatively neutral treatment of graduate students. Ministral is invariant in using “dear” for all groups, indicating no formality-based differentiation. The remaining models (Llama-3.1, Gemma-2, Claude-3.5) show moderate negative patterns with smaller coefficient magnitudes.
@@
For “research,” all models show positive associations with research-oriented patron types, but the magnitude varies considerably. Gemini-2.5 exhibits the strongest research accommodation with coefficients of +32.6 for faculty, +24.2 for outside users, and +17.6 for graduate students, indicating pronounced recognition of research hierarchies. Other models demonstrate more modest but consistent patterns, with faculty and outside users receiving positive associations across all LLMs, while graduate students receive variable treatment. Notably, treatment of staff members is inconsistent across models, ranging from negative associations in Llama-3.1 (--0.8) and GPT-4o (--2.4) to positive associations in others. This reflects LLMs’ uncertainty about staff research involvement, mirroring real-world variability in which some staff are research-intensive (e.g., systems programmers in medical schools) while others focus on transactional services (circulation desk staff or facilities management).
6 Discussion
6.1 Demographic Neutrality
@@
Our findings reveal encouraging evidence that contemporary LLMs largely avoid the demographic biases that have historically beset human-delivered reference services. Across race, ethnicity, and sex, we observed minimal systematic variation in LLM responses, a marked contrast to the substantial disparities documented in audit studies of human librarians (Shachaf and Horowitz,, 2006; Hamer,, 2021; Vladoiu et al.,, 2023). This neutrality represents a significant advancement over earlier language models that exhibited pronounced stereotypical biases in benchmark evaluations (Nadeem et al.,, 2021; Dhamala et al.,, 2021). The near-absence of racial and ethnic bias is particularly noteworthy given that LLMs are trained on vast text corpora that inevitably contain historical prejudices (Bender et al.,, 2021).
@@
These findings align with recent evaluations of contemporary models: OpenAI’s internal studies of GPT-4 found negligible differences in response quality by gender presentation and far less than 1% incidence of harmful stereotypes (Eloundou et al.,, 2025). However, improvements in bias reduction are not uniform across all models or contexts, as recent studies have documented varying degrees of bias mitigation across different architectures and training procedures (Gallegos et al.,, 2024). This underscores the importance of preemptive bias auditing, particularly in specialized domains such as academic libraries, where model-specific evaluation remains essential for ensuring equitable service delivery. While the lone significant result from Llama-3.1 warrants monitoring, the overall pattern suggests that recent advances in LLM safety (Ouyang et al.,, 2022; Bai et al.,, 2022) may effectively mitigate demographic bias in library applications.
@@
Llama-3.1’s increased use of “dear” for female patrons could reflect what Herring, (1994) describes as gendered politeness patterns in digital communication. This pattern may also stem from various sources: the model’s training data might contain more examples of “Dear [Name]” in contexts addressing women, reflecting historical norms of letter writing or email etiquette, or it could result from artifacts in the RLHF process wherein human raters inadvertently encoded gendered expectations about politeness. Rather than “deciding” to be more polite to women, the model could be reproducing subtle patterns learned during training. While this represents a relatively mild pattern rather than discriminatory content, it echoes early research on computer-mediated communication showing that women are associated with more elaborate positive politeness strategies in digital environments (Herring,, 1992). These patterns warrant investigation in future work to better understand the sources of such linguistic variations across different stages of model development.
6.2 Politeness Strategies & Formality Management
@@
The contrasting patterns of “dear” and “regards” usage in LLM responses reveal systematic formality strategies that merit interpretation through sociolinguistic and discourse-pragmatic frameworks. Drawing on Irvine’s formality framework (Irvine,, 1979), these patterns suggest “invoking positional identities,” whereby LLMs adjust language to reflect inferred relationships between patrons and institutions. From a politeness theory perspective, these patterns align with Brown and Levinson’s framework, which predicts more negative politeness strategies (e.g., deference, formality) in high-distance relationships and more positive strategies (e.g., warmth, solidarity) in low-distance ones (Brown and Levinson,, 1987). The preferential use of “dear” with graduate students, along with its avoidance for faculty, staff, alumni, and outside users, may well reflect learned associations regarding appropriate relational positioning in academic settings. Graduate students often occupy mentorship-receiving roles, where a warmer salutation connotes support. Conversely, omitting “dear” with higher-status or external patrons helps avoid face-threatening overfamiliarity or misaligned deference.
@@
The complementary use of “regards” as a formal closing—more common in faculty- and staff-oriented responses—functions as a distancing but respectful boundary marker. The observed pairing of “dear” (opening) and “regards” (closing) exemplifies what Irvine terms “code consistency,” wherein formality markers carry coherent social meanings across sequential positions in interaction. This aligns with the norms of institutional genre contexts, particularly academic emails, where role-indexed formality is conventional (Bazerman,, 1994; Bhatia,, 1993).
@@
Taken together, these usage patterns reveal the model’s capacity to mirror institutional facework and politeness strategies in context-sensitive ways. Future research should investigate whether these salutation and closing patterns persist across diverse prompt formulations and model architectures. Understanding whether such linguistic choices arise from training data biases or emergent communicative behavior remains essential to evaluating the fairness and appropriateness of LLM-mediated library interactions.
6.3 Communication Accommodation in Library Reference Service
@@
Patterns in LLM outputs suggest that models approximate institutional accommodation strategies, consistent with the convergence and divergence principles of CAT (Giles et al.,, 1991). These strategies manifest as variations in formality, professional self-references, and domain-specific vocabulary across different patron types to align with perceived roles, expertise levels, and institutional familiarity.
@@
One notable case of convergence appears in the increased use of “librarian” as a self-referential professional marker when responding to outside users and alumni. This aligns with institutional talk conventions, where explicit identity signaling is used to establish expertise and authority (Drew and Heritage,, 1992). It also reflects genre-based expectations in academic reference discourse, particularly for users less familiar with current institutional norms (Swales,, 1990). These identity markers may have been learned from training corpora containing email exchanges or reference interactions where such cues were more frequent in external communications.
@@
Similarly, LLMs use research-oriented vocabulary more frequently when responding to faculty, graduate students, and external researchers, suggesting that models adapt content to perceived domain expertise. This is consistent with CAT-informed audience design, wherein communicators tailor discourse to match assumed knowledge levels. Faculty and graduate students receive direct, specialized suggestions, while undergraduates and staff receive more generic or instructional responses, paralleling human librarian behavior.
@@
7 Conclusion
This study provides the first fairness evaluation of state-of-the-art LLMs deployed in virtual library reference services. We examined systematic differences in how LLMs respond to various user groups (with respect to sex, race/ethnicity, and patron type). We found LLMs do not introduce demographic bias when prompted to work as a “helpful, respectful, and honest librarian,” and they can approximate situational appropriateness by modulating tone, content, and stance in line with institutional norms. While human librarians remain essential for complex, ethical, or affective interactions, LLMs show promising readiness to offer equitable support for reference tasks.
Additionally, we contribute a protocol to detect and understand bias in LLM-based applications: the Fairness Evaluation Protocol. Library technologists can use this framework to audit the fairness of LLM-based library services and choose approaches that align with institutional principles and values (e.g., preferred degree of accommodation). This protocol is model-agnostic and generalizable to other domains adopting LLMs. It requires no access to model weights, making it well suited for practical evaluation and continuous monitoring by institutions concerned with LLM accountability.
@@ 7 Conclusion
Limitations & Future Directions
Several limitations and future directions are identified. First, our use of binary sex categories, while practically necessary for this initial investigation, does not capture the full spectrum of gender identities that libraries increasingly seek to serve. Second, our synthetic persona approach, while enabling controlled experimentation, may not fully capture the complexity of real-world patron interactions, including varied communication styles, cultural contexts, and intersectional identities that could influence LLM responses. Our study examined sex and race/ethnicity as separate dimensions, but real-world bias often operates at intersections (e.g., differential treatment of Black women versus White men), which warrants investigation in future work (Crenshaw,, 1989). Third, our findings are limited to formally written English queries. Recent research demonstrates that LLMs can exhibit context-dependent bias: minimal discrimination when demographic identity is implied through names, yet pronounced bias when users employ dialect variations (Hofmann et al.,, 2024). If input styles change (e.g., patrons writing in African American English or other dialects), biases that remained latent in our evaluation might emerge. Future work should address these limitations by adopting more inclusive demographic categories, examining intersectional effects, testing diverse linguistic styles, and systematically varying prompt designs to better understand their effects.
Despite these limitations, our FEP to detect systematic differences aligns with emerging best practices in LLM bias evaluation. Recent frameworks such as OpenAI’s bias enumeration tool (Eloundou et al.,, 2025) and LangFair (Bouchard,, 2025) similarly employ classifiers to identify output variations across demographic groups. This convergence across research and industry contexts reinforces the validity of using diagnostic classifiers for bias detection, particularly when combined with interpretive analysis to distinguish between harmful discrimination and appropriate contextual adaptation (Dwork et al.,, 2012; Binns,, 2020; Katell et al.,, 2020).
@@ Acknowledgments
Appendix A Dataset Balance Statistics
To ensure the validity of our fairness evaluation, we verified that our synthetic dataset maintains balanced representation across all demographic and institutional categories. We assessed class balance using the imbalance ratio metric, defined as the proportion of the largest class divided by the proportion of the smallest class within each demographic axis. An imbalance ratio of 1.0 indicates perfect balance, while higher values indicate increasing imbalance. Following best practices for machine learning fairness studies, we adopted a strict threshold of 1.5, below which distributions are considered well-balanced.
Table 6 presents balance statistics across the three demographic axes for all models. All models demonstrate excellent balance across sex and race/ethnicity categories, with imbalance ratios consistently below 1.06. Patron type distributions show slightly higher variation but remain well within acceptable bounds, with all ratios below 1.2.
@@ Appendix A Dataset Balance Statistics
Appendix B Classifier Hyperparameters
Our diagnostic classifiers are intentionally configured with conservative hyperparameters to ensure robust bias detection while avoiding overfitting to spurious patterns. Such conservative hyperparameter choices prioritize robustness over maximum performance, as our goal is bias detection rather than classification optimization. Table 7 presents the specific configurations for each classifier.
@@ Appendix B Classifier Hyperparameters
Logistic Regression C 1.0 (moderate regularization strength)
@@
Activation ReLU (standard nonlinear activation)
Solver Adam (adaptive learning rate)
Alpha 1e-4 (light L2 regularization)
@@
Column sample 0.8 (feature sampling for robustness)
Regularization $\alpha=0.1$, $\lambda=1.0$ (L1 and L2 penalties)
@@
Appendix C Fairness Evaluation Results
C.1 Race/Ethnicity Classification
@@ C.2 Sex Classification
C.3 Patron Type Classification
@@
Appendix D Radar Plot Coefficient Data

### v3

@@ Introduction
Beyond domain-specific vocabulary, LLMs also modulate service-oriented language based on assumptions about user autonomy and institutional familiarity. The use of “librarian” represents self-referential professional identification, as the LLM identifies itself as the service provider, and presents a particularly interesting case of professional authority assertion. This term appears more frequently in responses to outside users (+4.3) and alumni (+2.5), functioning as a form of institutional self-identification. This pattern may suggest that LLMs strategically deploy professional identity markers when interacting with patron types who are more likely to respond to explicit signals of institutional expertise and legitimacy as a means of establishing credibility (drew2006institutional).
The term “help” shows strong positive associations with outside users (+8.5) and faculty (+4.6), suggesting explicit assistance framing that serves different functions: providing institutional guidance for outside users while offering specialized research support for faculty. Conversely, “access” demonstrates consistent negative associations across most patron types, with particularly strong negative coefficients for graduate students (-8.1), alumni (-7.8), and outside users (-7.8). This pattern suggests that LLMs assume these groups either have existing access privileges or require different types of resource provision than undergraduates. The term “questions” shows negative associations with most patron types; these are particularly strong for graduate students (-11.1) and alumni (-9.1). This potentially indicates reduced meta-discourse about inquiry processes for users assumed to have greater institutional sophistication.
LLM-specific behaviours
@@
While the aggregate patterns demonstrate consistent role-based accommodation across LLMs, individual LLMs exhibit variable behaviours in implementing those accommodations. To illustrate this cross-model heterogeneity, we focus on three representative consensus features: “thank,” “dear,” and “research.” Figure 4 presents radar plots displaying each LLM’s coefficients for these features across patron types (see Supplementary Radar plot coefficients for details).
@@
The “thank” panel reveals substantial variation in accommodation strategies across models. Three models (Gemma-2, GPT-4o, and Claude-3.5) exhibit zero coefficients across all patron types, indicating uniform usage regardless of user identity. In contrast, two models implement pronounced hierarchical accommodation: Llama-3.1 demonstrates positive associations across all patron types (staff +39.3, faculty +31.8, graduate students +25.3), while Gemini-2.5 shows similar positive patterns with different magnitudes (graduate students +32.0, staff +28.9, alumni +21.3). Ministral exhibits a distinct pattern with consistent negative coefficients across all groups (-5.9 to -12.0), suggesting reduced usage of “thank” for non-undergraduate patrons.
@@
The “dear” panel illustrates substantial model-specific variation in formality management approaches. GPT-4o exhibits the most pronounced pattern: positive coefficients for graduate students (+53.6) and faculty (+44.9) but negative coefficients for staff (-28.7) and outside users (-35.9), implementing a selective graduate/faculty-focused formality strategy. Gemini-2.5 is most formal toward undergraduate students. It otherwise demonstrates consistent formality reduction with negative coefficients across most patron types (-51.8 to -60.9), except for relatively neutral treatment of graduate students. Ministral is invariant in using “dear” for all groups, indicating no formality-based differentiation. The remaining models (Llama-3.1, Gemma-2, Claude-3.5) show moderate negative patterns with smaller coefficient magnitudes.
@@
For “research,” all models show positive associations with research-oriented patron types, but the magnitude varies considerably. Gemini-2.5 exhibits the strongest research accommodation with coefficients of +32.6 for faculty, +24.2 for outside users, and +17.6 for graduate students, indicating pronounced recognition of research hierarchies. Other models demonstrate more modest but consistent patterns, with faculty and outside users receiving positive associations across all LLMs, while graduate students receive variable treatment. Notably, treatment of staff members is inconsistent across models, ranging from negative associations in Llama-3.1 (-0.8) and GPT-4o (-2.4) to positive associations in others. This reflects LLMs’ uncertainty about staff research involvement, mirroring real-world variability in which some staff are research-intensive (e.g., systems programmers in medical schools) while others focus on transactional services (circulation desk staff or facilities management).
Discussion
Demographic neutrality
@@ Discussion
Our findings reveal encouraging evidence that contemporary LLMs largely avoid the demographic biases that have historically beset human-delivered reference services. Across race, ethnicity, and sex, we observed minimal systematic variation in LLM responses, a marked contrast to the substantial disparities documented in audit studies of human librarians (shachaf2006virtual; hamer2021colour; vladoiu2023gender). This neutrality represents a significant advancement over earlier language models that exhibited pronounced stereotypical biases in benchmark evaluations (nadeem2021stereoset; dhamala2021bold). The near-absence of racial and ethnic bias is particularly noteworthy given that LLMs are trained on vast text corpora that inevitably contain historical prejudices (bender2021dangers).
@@
These findings align with recent evaluations of contemporary models: OpenAI’s internal studies of GPT-4 found negligible differences in response quality by gender presentation and an incidence of harmful stereotypes of far less than 1% (eloundou2025firstperson). However, improvements in bias reduction are not uniform across all models or contexts, as recent studies have documented varying degrees of bias mitigation across different architectures and training procedures (gallegos2024bias). This underscores the importance of pre-emptive bias auditing, particularly in specialized domains such as academic libraries, where model-specific evaluation remains essential for ensuring equitable service delivery. While the lone significant result from Llama-3.1 warrants monitoring, the overall pattern suggests that recent advances in LLM safety (ouyang2022training; bai2022constitutional) may effectively mitigate demographic bias in library applications.
@@
Llama-3.1’s increased use of “dear” for female patrons could reflect what herring1994politeness describes as gendered politeness patterns in digital communication. This pattern may also stem from various sources: the model’s training data might contain more examples of “Dear [Name]” in contexts addressing women, reflecting historical norms of letter writing or email etiquette, or it could result from artefacts in the RLHF process wherein human raters inadvertently encoded gendered expectations about politeness. Rather than “deciding” to be more polite to women, the model could be reproducing subtle patterns learned during training. While this represents a relatively mild pattern rather than discriminatory content, it echoes early research on computer-mediated communication showing that women are associated with more elaborate positive politeness strategies in digital environments (herring1992participation). These patterns warrant investigation in future work to better understand the sources of such linguistic variations across different stages of model development.
Politeness strategies and formality management
@@
The contrasting patterns of “dear” and “regards” usage in LLM responses reveal systematic formality strategies that merit interpretation through sociolinguistic and discourse-pragmatic frameworks. Drawing on Irvine’s formality framework (irvine1979formality), these patterns suggest “invoking positional identities,” whereby LLMs adjust language to reflect inferred relationships between patrons and institutions. From a politeness theory perspective, these patterns align with Brown and Levinson’s framework, which predicts more negative politeness strategies (e.g., deference, formality) in high-distance relationships and more positive strategies (e.g., warmth, solidarity) in low-distance ones (brown1987politeness). The preferential use of “dear” with graduate students, along with its avoidance for faculty, staff, alumni, and outside users, may well reflect learned associations regarding appropriate relational positioning in academic settings. Graduate students often occupy mentorship-receiving roles, where a warmer salutation connotes support. Conversely, omitting “dear” with higher-status or external patrons helps avoid face-threatening overfamiliarity or misaligned deference.
@@
The complementary use of “regards” as a formal closing—more common in faculty- and staff-oriented responses—functions as a distancing but respectful boundary marker. The observed pairing of “dear” (opening) and “regards” (closing) exemplifies what Irvine terms “code consistency,” wherein formality markers carry coherent social meanings across sequential positions in interaction. This aligns with the norms of institutional genre contexts, particularly academic emails, where role-indexed formality is conventional (bazerman1994systems; bhatia1993analyzing).
@@
Taken together, these usage patterns reveal the model’s capacity to mirror institutional facework and politeness strategies in context-sensitive ways. Future research should investigate whether these salutation and closing patterns persist across diverse prompt formulations and model architectures. Understanding whether such linguistic choices arise from training data biases or emergent communicative behaviour remains essential to evaluating the fairness and appropriateness of LLM-mediated library interactions.
Communication accommodation in library reference service
@@
Patterns in LLM outputs suggest that models approximate institutional accommodation strategies, consistent with the convergence and divergence principles of CAT (giles1991accommodation). These strategies manifest as variations in formality, professional self-references, and domain-specific vocabulary across different patron types to align with perceived roles, expertise levels, and institutional familiarity.
@@
One notable case of convergence appears in the increased use of “librarian” as a self-referential professional marker when responding to outside users and alumni. This aligns with institutional talk conventions, where explicit identity signalling is used to establish expertise and authority (drew2006institutional). It also reflects genre-based expectations in academic reference discourse, particularly for users less familiar with current institutional norms (swales1990genre). These identity markers may have been learned from training corpora containing email exchanges or reference interactions where such cues were more frequent in external communications.
@@
Similarly, LLMs use research-oriented vocabulary more frequently when responding to faculty, graduate students, and external researchers, suggesting that models adapt content to perceived domain expertise. This is consistent with CAT-informed audience design, wherein communicators tailor discourse to match assumed knowledge levels. Faculty and graduate students receive direct, specialized suggestions, while undergraduates and staff receive more generic or instructional responses, paralleling human librarian behaviour.
@@
Conclusion
This study provides the first fairness evaluation of state-of-the-art LLMs deployed in virtual library reference services. We examined systematic differences in how LLMs respond to various user groups (with respect to sex, race/ethnicity, and patron type). We found LLMs did not introduce demographic bias when prompted to work as a “helpful, respectful, and honest librarian,” and they can approximate situational appropriateness by modulating tone, content, and stance in line with institutional norms. While human librarians remain essential for complex, ethical, or affective interactions, LLMs show promising readiness to offer equitable support for reference tasks.
Additionally, we contribute a protocol to detect and understand bias in LLM-based applications: the Fairness Evaluation Protocol. Library technologists can use this framework to audit the fairness of LLM-based library services and choose approaches that align with institutional principles and values (e.g., preferred degree of accommodation). This protocol is model-agnostic and generalisable to other domains adopting LLMs. It requires no access to model weights, making it well-suited for practical evaluation and continuous monitoring by institutions concerned with LLM accountability.
@@ Conclusion
Limitations and future directions
Several limitations and future directions are identified. First, our use of binary sex categories, while practically necessary for this initial investigation, does not capture the full spectrum of gender identities that libraries increasingly seek to serve. Second, our synthetic persona approach, while enabling controlled experimentation, may not fully capture the complexity of real-world patron interactions, including varied communication styles, cultural contexts, and intersectional identities that could influence LLM responses. Our study examined sex and race/ethnicity as separate dimensions, but real-world bias often operates at intersections (e.g., differential treatment of Black women versus White men), which warrants investigation in future work (crenshaw1989intersectionality). Third, our findings are limited to formally written English queries. Recent research demonstrates that LLMs can exhibit context-dependent bias: minimal discrimination when demographic identity is implied through names, yet pronounced bias when users employ dialect variations (hofmann2024ai). If input styles change (e.g., patrons writing in African American English or other dialects), biases that remained latent in our evaluation might emerge. Future work should address these limitations by adopting more inclusive demographic categories, examining intersectional effects, testing diverse linguistic styles, and systematically varying prompt designs to better understand their effects.
Institutions adopting FEP should involve librarians throughout the audit process. FEP detects whether responses differ across groups, but librarians must interpret whether detected patterns constitute appropriate accommodation or problematic bias. Librarians are best positioned to refine system prompts, audit for factual errors, and determine acceptable response styles given institutional priorities and resource constraints. Beyond computational auditing, patron perception surveys and ethnographic observation of pilot deployments can reveal service disparities that controlled experiments miss. Such triangulation across computational analysis, professional judgment, and lived experience can provide a full view of LLM service fairness.
Despite these limitations, our FEP to detect systematic differences aligns with emerging best practices in LLM bias evaluation. Recent frameworks such as OpenAI’s bias enumeration tool (eloundou2025firstperson) and LangFair (bouchard2025langfair) similarly employ classifiers to identify output variations across demographic groups. This convergence across research and industry contexts reinforces the validity of using diagnostic classifiers for bias detection, particularly when combined with interpretive analysis to distinguish between harmful discrimination and appropriate contextual adaptation (dwork2012fairness; binns2020apparent; katell2020toward).
@@ Acknowledgments
Dataset balance statistics
To ensure the validity of our fairness evaluation, we verified that our synthetic dataset maintains balanced representation across all demographic and institutional categories. We assessed class balance using the imbalance ratio metric, defined as the proportion of the largest class to the proportion of the smallest class within each demographic axis. An imbalance ratio of 1.0 indicates perfect balance, while higher values indicate increasing imbalance. Following best practices for machine learning fairness studies, we adopted a strict threshold of 1.5, below which distributions are considered well-balanced.
Table 6 presents balance statistics across the three demographic axes for all models. All models demonstrate excellent balance across the sex and race/ethnicity categories, with imbalance ratios consistently below 1.06. Patron type distributions show slightly greater variation but remain well within acceptable bounds, with all ratios below 1.2.
@@
Classifier hyper-parameters
Our diagnostic classifiers are intentionally configured with conservative hyper-parameters to ensure robust bias detection while avoiding over-fitting to spurious patterns. Such conservative hyperparameter choices prioritize robustness over maximum performance, as our goal is bias detection rather than classification optimization. Table 7 presents the specific configurations for each classifier.
@@
Logistic Regression C 1.0 (moderate regularisation strength)
@@
Activation ReLU (standard non-linear activation)
Solver Adam (adaptive-learning optimiser)
Alpha 1e-4 (light L2 regularisation)
@@
Column sample 0.8 (feature-sampling for robustness)
Regularisation $\alpha=0.1$, $\lambda=1.0$ (L1 and L2 penalties)
@@
Fairness evaluation results
Race/Ethnicity classification
@@
Patron type classification
@@
Llama-3.1 temperature sensitivity analysis
We conducted experiments with Llama-3.1 8B at temperatures 0.0 (greedy decoding producing deterministic output) and 0.3 (low variance), comparing results with our primary analysis at temperature 0.7.
We generated 2,500 responses per temperature setting across five random seeds. At temperature 0.0, we observed minimal duplication (4 duplicates across 2,500 responses). Average response lengths were comparable across temperatures: 222 words at $T=0.7$ [218, 226], 229 words at $T=0.0$ [223, 235], and 228 words at $T=0.3$ [223, 234]. We applied identical analysis pipelines with Bonferroni-corrected significance testing ($\alpha=0.05/6=0.0083$, accounting for three classifiers × two temperature conditions per characteristic).
Fairness patterns remain qualitatively consistent across temperatures, confirming that our conclusions are robust to the choice of sampling temperature. For sex, the same linguistic feature (“dear”) drives classification across all temperatures. At $T=0.7$, only logistic regression reached significance (+5.00%*), whereas at $T=0.0$ and $T=0.3$, all three classifiers achieved significance (Table 12), with XGBoost showing the strongest performance (+8.00%* and +8.28%* respectively). Volcano plot analysis (Figure 5) confirms that “dear” remains the sole statistically significant feature at $T=0.0$ and $T=0.3$, replicating the pattern observed at $T=0.7$. This consistency demonstrates that the sex-related pattern reflects a genuine systematic preference in salutation style rather than temperature-dependent artefact.
(a) Temp = 0.0
(b) Temp = 0.3
Figure 5: Volcano plots for sex classification at temperatures 0.0 and 0.3. Each plot shows feature importance (x-axis: log-odds coefficients) versus statistical significance (y-axis: $\log_{10}(p)$). The salutation “dear” remains the only statistically significant feature after correction across both temperature settings, consistent with findings at $T=0.7$.
For race/ethnicity, the marginally significant pattern observed at $T=0.7$ (LogReg +1.93%* margin) disappeared entirely at both $T=0.0$ (+1.01%) and $T=0.3$ (-0.19%), with no classifiers reaching significance (Table 11). This absence of replication at lower temperatures confirms that the $T=0.7$ result represented a marginal statistical fluctuation rather than genuine racial/ethnic bias, strengthening our conclusion that Llama-3.1 maintains demographic neutrality across race and ethnicity groups.
Patron type classification remained significant across all temperatures (Table 13), exhibiting the same pattern of institutional role accommodation with amplified effect sizes at lower temperatures: $T=0.0$ showed margins of +20.69%* to +35.49%*, compared to +15.13%* to +19.21%* at $T=0.3$ and +10.05%* to +11.05%* at $T=0.7$. The consistent significance across all temperature settings confirms that institutional role accommodation is not sensitive to temperature choice.
Table 11: Classification performance margins over random chance (16.67%) for predicting user race/ethnicity from Llama-3.1 responses. Asterisks (*) indicate significance after Bonferroni correction ($\alpha=0.0083$ for ablation; $\alpha=0.0028$ for primary analysis).
Temperature LogReg MLP XGBoost
0.0 +1.01 +0.21 +1.01
0.3 -0.19 -1.19 -0.03
0.7 (primary) +1.93* +0.97 +1.61
Table 12: Classification performance margins over random chance (50.00%) for predicting user sex from Llama-3.1 responses. Asterisks (*) indicate significance after Bonferroni correction ($\alpha=0.0083$ for ablation; $\alpha=0.0028$ for primary analysis).
Temperature LogReg MLP XGBoost
0.0 +5.52* +4.68* +8.00*
0.3 +4.96* +3.68* +8.28*
0.7 (primary) +5.00* +1.96 +2.32
Table 13: Classification performance margins over random chance (16.67%) for predicting user patron type from Llama-3.1 responses. All results significant after Bonferroni correction ($\alpha=0.0083$ for ablation; $\alpha=0.0028$ for primary analysis).
Temperature LogReg MLP XGBoost
0.0 +20.69* +33.77* +35.49*
0.3 +15.13* +15.29* +19.21*
0.7 (primary) +10.05* +10.73* +11.05*
Radar plot coefficients